if i traveled 3 hours at an average rate of 60mph. due to traffic it took me 5 hours hours at an average rate of 36mph. what is my average of my entire trip
Answer by njusa(12)

60*3 + 5*36 = 360.... total distance
3 + 5 = 8 ..... total time
360/8 = 45.. average speed for the trip.
Mukhi